what is the domain and range of the inverse of the f(x)= Sqrt(x-4)
Answer by edjones(8007)

f(x)= Sqrt(x-4)
y=sqrt(x-4)
x=sqrt(y-4)
x^2=y-4
y=x^2+4
.
The domain is all real numbers.
The range is 4 to infinity.
.
Ed